OpenStudy (anonymous):

If f(x) = 2x^2 + 3x and g(x) = x – 2, what is (f + g)(2)?

OpenStudy (anonymous):

A)16 B)14 C)12 D)10 E)8

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle f(x) = 2x^2 + 3x }\) \(\large\color{green}{ \displaystyle g(x) = x – 2 }\) \(\large\color{black}{ \displaystyle \color{blue}{f(x) }+ \color{green}{g(x)} ~~~~\Longrightarrow ~~~~( \color{blue}{2x^2+3x})~+~(\color{green}{x-2})}\)